OpenStudy (anonymous):

The center of a circle is at (4,-1) and it has a radius of 6. What is the equation of the circle? (Points : 3)

Directrix (directrix):

Write the equation in the format shown below on the attachment.

OpenStudy (anonymous):

Ok...

Directrix (directrix):

@PamT And, what did you get as an answer?

OpenStudy (anonymous):

I picked a @Directrix that's wrong?

OpenStudy (anonymous):

(x 4)2 + (y + 1)2 = 6 (x 4)2 + (y + 1)2 = 36 (x + 4)2 + (y 1)2 = 6 (x + 4)2 + (y 1)2 = 36 those r the choices

Directrix (directrix):

@PamT >> picked a that's wrong? You forgot to square the radius. (x - 4)^2 + (y - (-1))^2 = 6^2 Which option is this ^^^